A ball is projected vertically upwards with a velocity of 24.5m/s from the bottom of the tower.A boy who is standing at the top of the tower is unable to catch the ball when it passes him in the upward direction. But the ball again reaches after 3sec,when it is falling and then he catches it. What is the height of the tower?

Grade:11

1 Answers

Lok Nitesh
5 years ago
I think so the height of the tower is 62.5 
meter as t = h/u 
h=u×t=24.5 × 3= 62.5 .                                                                                                                          I don't whether it is correct so plz check it